Economic dispatch. The fuel cost function in dollars / hour of two
thermal plants is
C1=320+6.2 P1+0.004 P1^2
C2=200+6.0 P2+0.003 P2^2
where P1 and P2 are in MW. Plant outputs are subject to the following constraints:
50 ≤ P1 ≤ 250
50 ≤ P2 ≤ 350
On a 100MVA base, the per unit real power loss is
PL = 0.0125 P1^2+ 0.00625 P2^2
The total load is 412.35MW. Determine the optimal dispatch of generation.
Answer #1

☺ - 320+6.1P, +0.004842 C2 = 200 + 6-08, +0.003Pi PL :0.01257, +0.00625 P? 50 sp, ¢ 250 - 50 € P2 & 350 Power Demand= 412.356.1.-6.0 -0.076251,-0.006Pqf0.008P+0:15 Pi -0.00012, P+0.00015P, P2 : 0 - © -0.08225 P2 +0:1588 to 00005 P, PR-0 Also Pitp2 =